# FusionSAM: Latent Space driven Segment Anything Model for Multimodal Fusion and Segmentation

### v1

@@ Introduction
$\left\{Q_{0},K_{0},V_{0}\right\}=\left\{z_{0}^{q}W_{Q1},z_{0}^{q}W_{K1},z_{0}^{q}W_{V1}\right\},$ (12)
$\left\{Q_{f},K_{f},V_{f}\right\}=\left\{z_{f}^{q}W_{Q2},z_{f}^{q}W_{K2},z_{f}^{q}W_{V2}\right\},$
$z_{f}$ $=\text{LN}\left(\text{softmax}\left(\frac{Q_{f}K_{o}^{T}}{\sqrt{d_{k}}}\right)V_{o}\right)+z_{c}.$ (13)
The final representation $z_{f}$ is then processed through a convolutional layer to produce the fusion mask $I_{f}$, which serves as a precise prompt for guiding pixel-level segmentation.
@@
Figure 4: Visualization of ablation studies in FusionSAM.
Method mIoU(%)
MFNet FMB
(A) 35.6 41.4
(B) 47.3 57.6
(C) FusionSAM 63.0 61.8
Table 2: Ablation study results for FusionSAM.
@@
Figure 5: Qualitative demonstrations of different approaches on the FMB dataset.
@@
Ours 80.1 52.8 45.9 43.7 46.4 85.5 88.4 50.8 61.8
Table 3: Results of quantitative segmentation on the test set of FMB dataset.
Implementation Details
Datasets
@@
Two representative datasets, including MFNet (Ha et al. 2017) and FMB (Liu et al. 2023), containing 1569 and 1500 pairs of visible and infrared images with resolutions of 480$\times$640 and 600$\times$800, respectively, to train and evaluate our method. Annotated into 9 and 14 categories relevant to autonomous driving and semantic understanding, these datasets offer varied lighting conditions and rich scenes that enhance the generalization ability of fusion and segmentation models.
Training Detail
@@
During 100 epochs of training, multimodal images are subjected to 4$\times$ downsampled features by the LSTG module, and the FMP module further captures efficient fusion representations, combined with 10-point mask prompts and 1-box mask prompt to facilitate effective segmentation of SAM. Our initial learning rate is set to 1e-4, using the Adam optimizer with a weight decay of 1e-3, the batch size is set to 4, and vit/h is used as the encoder. All experiments are performed on a NVIDIA A100 Tensor Core GPU. We use mean intersection over union (mIoU) to quantitatively evaluate the performance of semantic segmentation. mIoU is the average result of summing the ratio of the intersection over the sum of the predicted true values for each class.
Experimental Results
Ablation Study
To explore the contribution of each part of our method in detail, we designed three scenarios: (A) omitting the LSTG module compared to our FusionSAM; (B) removing the FMP module from the fusion process and replacing it with direct concat; (C) complete FusionSAM. The results of the ablation experiment are shown in Figure 4. We can observe that FusionSAM achieves the best segmentation results on both datasets. As shown in Table 4, in (A) , by removing the LSTG module, we notice that the results drop by 27.4% and 20.4%, respectively, while resulting in poor segmentation results, which shows the effectiveness of the LSTG module in generating latent space tokens through vector quantization. Our fusion method is verified in (B). Without introducing the fusion mask hint, the model has difficulty distinguishing the foreground and background, ignoring the unique and complementary features of each modality, resulting in a decrease in mIoU of 15.7% and 4.2%, respectively. Therefore, our proposed LSTG module and FMP module can effectively improve the segmentation performance of multimodal images and produce excellent visual results.
Comparisons with Previous Methods
Comparison with SAM.
@@
SAM (Kirillov et al. 2023) is competitive in the segmentation field because of its powerful segmentation performance and adaptability in different fields. Compared with SAM, SAM2 (Ravi et al. 2024) has significant improvements in applicable fields, segmentation accuracy, and running speed. To demonstrate the effective design and powerful performance of our FusionSAM and maintain a fair comparison, we use SAM and SAM2 to directly infer the fused feature maps generated in FusionSAM, and the results are shown in Table 4. The SAM series cannot handle multimodal image segmentation, whereas our method introduces SAM into the multimodal field, ensuring its excellent segmentation performance and expanding its applicability in more complex scenarios.
Method mIoU(%)
MFNet FMB
SAM 32.7 34.6
SAM2 43.0 46.3
FusionSAM 63.0 61.8
Table 4: Performance comparison on the two datasets.
Comparison with SOTA.
@@
We conduct comparative experiments and evaluations with seven state-of-the-art semantic segmentation methods, including EGFNet (Zhou et al. 2022), SegMiF (Liu et al. 2023), EAEFNet (Liang et al. 2023), LASNet (Li et al. 2022), SFAF-MA (He et al. 2023), ECFNet (Shen et al. 2024), and MRFS (Zhang et al. 2024). We provide quantitative results in Tables 1 and 3. Our FusionSAM achieves the highest mIoU on both datasets. Compared with the second-highest method, FusionSAM improves mIoU by 3.9% and 0.6% on MFNet and FMB, respectively. More specifically, for heat-insensitive categories, such as Car Stop, Building, Curve, and Bump, our method achieves significant superiority due to the effective visual quality preservation and enhancement. Overall, these findings confirm that our method achieves SOTA excellence in semantic segmentation.
Result Visualization
@@
Figure 3 and Figure 5 show the visualization results on the MFNet and FMB datasets, compared with the most competitive methods. These datasets present segmentation challenges due to their rich categories, complex imaging conditions, and diverse scene details. Existing fusion methods struggle to highlight dim infrared targets (e.g., bicycles in Figure 3, second row) and recognize distant pedestrians (Figure 5, third row). Methods relying on two-stream networks often introduce conflicts if feature fusion is incomplete, leading to misclassifications, such as occluded cars (Figure 3, first row) and human shapes (Figure 5, first row). Additionally, edge blurring in dense target predictions is common (Figure 3, third row). By embedding latent space representations and achieving cross-modal consistency, our method reduces redundancy while retaining key information, significantly improving SAM’s segmentation performance and enabling accurate object classification across diverse scenes.
Conclusion
A key challenge in multimodal semantic segmentation for autonomous driving is developing a framework that can effectively fuse and utilize multimodal data as prompts during training, guiding the model to achieve high-performance segmentation in dense distribution scenes--- -an issue that previous multimodal segmentation approaches have not fully addressed. We have innovatively proposed FusionSAM, a latent space driven SAM framework for multimodal semantic segmentation, which endows the SAM architecture with robust capabilities in multimodal fusion, understanding, and segmentation. Our approach performs comprehensive cross-domain fusion of the latent space representations from two modalities, using this fused information as prompts to guide segmentation. This is the first study to leverage SAM in multimodal semantic segmentation of natural scenes, utilizing fusion as a guiding prompt. Extensive experiments demonstrate that FusionSAM significantly outperforms existing state-of-the-art methods in multimodal autonomous driving scenarios, offering a novel approach for future multimodal semantic segmentation tasks.

### v2

@@ 3.2.3. Fusion Mask Prompting Module
(11) $\left\{Q_{0},K_{0},V_{0}\right\}=\left\{z_{0}^{q}W_{Q1},z_{0}^{q}W_{K1},z_{0}^{q}W_{V1}\right\},$
$\left\{Q_{f},K_{f},V_{f}\right\}=\left\{z_{f}^{q}W_{Q2},z_{f}^{q}W_{K2},z_{f}^{q}W_{V2}\right\}.$
(12) $I_{f}$ $=\text{LN}\left(\text{softmax}\left(\frac{Q_{f}K_{o}^{T}}{\sqrt{d_{k}}}\right)V_{o}\right)+z_{c}.$
Equations 11 and 12 present the process of the Cross Convolutional layer of Figure 2. $z_{f}^{q}$ represents the result of directly performing Fusion Convolutional layer on the original latent space embeddings of the two modalities. The final representation is then processed through a convolutional layer to produce the fusion mask $I_{f}$, which serves as a precise prompt for guiding pixel-level segmentation. In previous SAM models, prompts were typically embedded as masks in the form of points or bounding boxes to guide segmentation. However, in FusionSAM, we innovate by using fused multimodal images as prompts. The ”fusion mask” refers to this integrated representation of both modalities, allowing the model to leverage comprehensive feature information for more accurate segmentation. This approach enhances the model’s ability to handle complex multimodal environments.
@@
Figure 4. A visual demonstration of different methods on the FMB dataset.
Table 2. Results of quantitative segmentation on the test set of FMB dataset.
@@
Ours 80.4 59.4 48.2 47.3 52.1 85.7 87.8 52.0 64.1
4. Experiment
4.1. Experimental Setups
@@
Datasets. Two representative datasets, including MFNet (Ha et al., 2017) and FMB (Liu et al., 2023), containing 1569 and 1500 pairs of visible and infrared images with resolutions of 480$\times$640 and 600$\times$800, respectively, to train and evaluate our method. Annotated into 9 and 14 categories relevant to autonomous driving and semantic understanding, these datasets offer varied lighting conditions and rich scenes that enhance the generalization ability of fusion and segmentation models. In addition, we select the VDT-2048 dataset, which consists of 2,048 pairs of RGB, thermal (T), and depth (D) images, spanning three distinct modalities. This dataset provides a comprehensive collection of multimodal data, designed to test and evaluate the performance of segmentation models across varying environmental conditions.
Figure 5. Qualitative demonstrations of different approaches on the VDT-2048 dataset. FusionSAM displays an RGB+D+T visualization. Our method shows optimal results in terms of texture details.
Table 3. Comparison results of $S,F_{\beta}^{\max},F_{\beta}^{\text{mean }},F_{\beta}^{adp},E_{\xi}^{\max},E_{\xi}^{\text{mean }},E_{\xi}^{adp}$ and $MAE$ on VDT-2048 for RGB-T SOD task. ” $\uparrow$ ” (” $\downarrow$ ”) means that the larger (smaller) the better. The best three results in each row are marked in red, green,The table contains a variety of visual modes including RGB+Depth, RGB+Thermal, and the contrast accuracy of RGB+Depth+Thermal.
Methods Modality $S$ $MAE$ $E_{\xi}^{adp}$ $E_{\xi}^{mean}$ $E_{\xi}^{max}$ $F_{\beta}^{adp}$ $F_{\beta}^{mean}$ $F_{\beta}^{max}$
BBSNet21(Zhai et al., 2021) RGB+D 0.9117 0.0046 0.8747 0.9357 0.9769 0.6957 0.8267 0.8679
Ours RGB+D 0.9407 0.0036 0.63 0.9512 0.9841 0.8707 0.8767 0.8912
DCNet22(Tu et al., 2022) RGB+T 0.8787 0.0038 0.9658 0.9436 0.9659 0.8521 0.8450 0.8525
LSNet23(Zhou et al., 2023) RGB+T 0.8867 0.0044 0.9327 0.9631 0.9737 0.7607 0.8097 0.8354
Ours RGB+T 0.9496 0.0029 0.9705 0.9790 0.9850 0.8822 0.8947 0.9056
HWSI22(Song et al., 2022) RGB+D+T 0.9318 0.0026 0.9815 0.9845 0.9890 0.8718 0.8961 0.9102
MFFNet23(Wan et al., 2023) RGB+D+T 0.9394 0.0025 0.9831 0.9825 0.9905 0.8758 0.9034 0.9180
Ours RGB+D+T 0.9542 0.0020 0.9890 0.9875 0.9952 0.8982 0.9108 0.9266
@@
Training Details. During 100 epochs of training, multimodal images are subjected to 4$\times$ downsampled features by the LSTG module, and the FMP module further captures efficient fusion representations. Our initial learning rate is set to 1e-4, using the Adam optimizer with a weight decay of 1e-3, the batch size is set to 4, and vit/h is used as the encoder. All experiments are performed on a NVIDIA A100 Tensor Core GPU. We use mean Intersection over Union (mIoU) to evaluate semantic segmentation performance. mIoU is the average of intersection over the union for each class.
4.2. Comparison with SAM
@@
SAM (Kirillov et al., 2023) is competitive in the segmentation field because of its powerful segmentation performance and adaptability in different fields. Compared with SAM, SAM2 (Ravi et al., 2024) has significant improvements in applicable fields, segmentation accuracy, and running speed. To demonstrate the effective design and powerful performance of our FusionSAM and maintain a fair comparison, we use SAM and SAM2 to directly infer the fused feature maps generated in FusionSAM, and the results are shown in Figure 7.(a). The SAM series cannot handle multimodal image segmentation, whereas our method introduces SAM into the multimodal field, ensuring its excellent segmentation performance and expanding its applicability in more complex scenarios.
4.3. Comparison with State-of-the-Arts
@@
We conduct comparative experiments and evaluations with seven state-of-the-art semantic segmentation methods, including EGFNet (Zhou et al., 2022), SegMiF (Liu et al., 2023), EAEFNet (Liang et al., 2023), LASNet (Li et al., 2022), SFAF-MA (He et al., 2023), ECFNet (Shen et al., 2024b), and MRFS (Zhang et al., 2024b). We provide quantitative results in Tables 1 and 2. Our FusionSAM achieves the highest mIoU on both datasets. Compared with the second-highest method, FusionSAM improves mIoU by 5.3% and 2.9% on MFNet and FMB, respectively. More specifically, for heat-insensitive categories, such as Bike, Car Stop, Building, Curve, and Truck, our method achieves significant superiority due to the effective visual quality preservation and enhancement. Overall, these findings confirm that our method achieves SOTA excellence in semantic segmentation.
As shown in Figure 5 and Table 3, we further demonstrate the generalization capability of FusionSAM by evaluating it on the RGB+Depth and RGB+Thermal modalities, as well as the combined RGB+Depth+Thermal modality. Our experiments show that FusionSAM consistently outperforms existing SOTA methods across all modalities. The fusion of all three modalities is performed using cross-attention to model the long-range dependencies between the features, which allows for the effective integration of complementary information from the different modalities. These results further highlight the robustness and effectiveness of our method across multiple input configurations, reinforcing its superiority in multimodal semantic segmentation tasks.
4.4. Ablation Study
Figure 6. Visualization of ablation studies in FusionSAM.
Figure 7. (a) Comparison of FusionSAM’s performance with the existing baseline models SAM and SAM2; (b) Performance comparison of FusionSAM in ablation experiments.
To explore the contribution of each part of our method in detail, we designed three scenarios: (A) Omitting the LSTG module compared to our FusionSAM; (B) Removing the FMP module from the fusion process and replacing it with direct concat; (C) Complete FusionSAM. The results of the ablation experiment are shown in Figure 6. We can observe that FusionSAM achieves the best segmentation results on both datasets. As shown in Figure 7.(b), in (A) , by removing the LSTG module, we notice that the results drop by 28.8% and 22.7%, respectively, while resulting in poor segmentation results, which shows the effectiveness of the LSTG module in generating latent space tokens through vector quantization. Our fusion method is verified in (B). Without introducing the fusion mask hint, the model has difficulty distinguishing the foreground and background, ignoring the unique and complementary features of each modality, resulting in a decrease in mIoU of 17.1% and 6.5%, respectively. Therefore, our proposed LSTG and FMP module can effectively improve the segmentation performance of multimodal images and produce excellent visual results.
4.5. Result Visualization
@@
Figures 3 and 4 show segmentation visualizations and Class Activation Mapping (CAM) of our method on the MFNet and FMB datasets, and compare with the most competitive methods. These datasets present segmentation challenges due to their rich categories, complex imaging conditions, and diverse scene details. Existing fusion methods struggle to highlight dim infrared targets (e.g., bicycles in Figure 3, second row) and recognize distant pedestrians (Figure 4, third row). Methods relying on two-stream networks often introduce conflicts if feature fusion is incomplete, leading to misclassifications, such as occluded cars (Figure 3, first row) and human shapes (Figure 4, first row). Additionally, edge blurring in dense target predictions is common (Figure 3, third row). By embedding latent space representations and achieving cross-modal consistency, our method reduces redundancy while retaining key information, significantly improving SAM’s segmentation performance and enabling accurate object classification across diverse scenes.
5. Conclusion
A key challenge in multimodal semantic segmentation for autonomous driving is efficiently fusing multimodal data as prompts to guide high-performance segmentation in dense scenes. Traditional methods often lead to redundancy and weakened modality information. To address this, we propose FusionSAM, a latent space-driven SAM framework that enhances multimodal fusion and segmentation. By utilizing vector quantization, we achieve non-redundant, efficient fusion while preserving modality-specific information. This is the first study to apply SAM to multimodal segmentation of natural scenes with fusion as a guiding prompt. Extensive experiments demonstrate that FusionSAM significantly outperforms state-of-the-art methods across multimodal visual scenes (RGB, infrared, depth), showing robustness in varying lighting conditions and offering a novel solution for future tasks.
Acknowledgements.
This work was supported in part by National Natural Science Foundation of China under Grant 62425109, 62322117, 62371365, U24B20136, and U22B2014, and the Fundamental Research Funds for the Central Universities under Grant ZYTS24101. This work was supported by the Fundamental Research Funds for the Central Universities, and supported by the Innovation Fund of Xidian University under Grant YJSJ25007. This work was supported in part by the Australian Research Council under Projects DP240101848 and FT230100549.